# Pulling apart the mechanisms that lead to jammed knitted fabrics

### v1

@@ Abstract
We investigate the mechanical behavior of jammed knitted fabrics, where geometric confinement leads to an initially stiff mechanical response that softens into low stiffness behavior with additional applied stress. We show that the jammed regime is distinguished by changes in yarn geometry and contact rearrangement that must occur to allow the individual stitches to stretch. These rearrangements allow for the relaxation of high residual contact forces that are present within jammed fabric, altering the low-strain response. We demonstrate that fabric jamming can be induced or reduced by changing either the constituent yarn or the fabric manufacturing parameters. Analysis of experimental samples shows that changing the “stitch size” in manufacturing affects both the in situ yarn radius and the length of yarn per stitch, both of which affect jamming.
@@ I Introduction
Granular systems are ubiquitous in nature and manufacturing, from sandy beaches to corn-filled silos, and they exhibit some common behaviors. Generally, jamming describes an amorphous system that is both disordered and has a yield stress [1, 2]. This disorder differentiates jamming from crystallization, which is both ordered and an equilibrium state. For athermal spheres, the onset of jamming signifies the sudden transition from zero to non-zero pressure response when the granular system is compressed [2]. A universal feature of granular systems is a jamming transition, where a system which has fluid-like behavior – flowing under constant stress – develops compressional rigidity and behaves like a solid under constant stress.
@@
The transition to or from a jammed state often occurs due to changes in constraints on the constituent particles. For hard particles, these constraints arise from contacts between neighbors, and the jamming transition can be characterized by the average number of contacts per particle [3]. There has been recent interest in “elasto-granular” systems that combine hard particle jamming with fiber elasticity to give rise to new mechanical behaviors [4, 5]. For moderately deformable systems, like foams, the number of contacts still plays a large role in the characterization of the jamming transition [6, 7]. In highly deformable systems – where the deformation is of the order of the particle size – additional constraints, such as particle shape [8, 9], contribute to the onset of jammed behavior. Jamming has been introduced as a particularly useful property in mechanical metamaterials and soft robotics, where the variable stiffness of the material leads to novel applications [10, 11, 12, 13]. In this vein, we examine jamming in a non-granular context, where the continuous, flexible yarn that is initially confined in tightly-knitted fabrics must rearrange in order to support higher stresses.
@@
Knitting is a fabric manufacturing method that creates some of the most commonly used clothing items, such as t-shirts and socks. Knitted fabrics are comprised of a network of interlocked loops formed by a continuous piece of yarn (Fig. 1a). In weft knitting, a single stand of yarn forms a series of loops horizontally within a row. In the next row, a new series of loops are threaded through the loops on the prior row. Each loop of yarn is called a stitch, and combining stitches with different topologies leads to different elastic responses and textures in the final fabric [14, 15]. A fabric made entirely of knit stitches is called stockinette fabric, or jersey, and is the fabric type found in the body of t-shirts. This is the type of fabric we will study here, and it is the stiffest of the four most common knitted fabric types [15, 16].
@@
The term “jamming”, as applied to knitted textiles, was coined in  [17] by Postle to describe an anomalous stiff response to low strains. Jamming in knitted fabrics is qualitatively different from jamming in granular systems in a multitude of ways. For one, knitted fabrics are far from amorphous, typically possessing periodic translational symmetry. Unlike jammed, athermal granular systems, the fabrics show jammed behavior under tension and move from a jammed state to a softened state (Fig. 1) with greater applied force [17]. Here, we argue that the change in mechanical properties due to yarn confinement in tightly-knitted fabrics can be regarded as analogous to the jamming behavior seen in other athermal, granular systems.
@@
Figure 1: (a) Portrayal of the simulated knitted fabric with a black box indicating the unit cell. This graphic is made from a simulation where $k=0.1$ $mN/mm^{-2}$ and $L=13.0$ $mm$ with an artificially small yarn radius to better display the yarn structure. (b) Simulated stress versus strain in the x-direction for two different fabrics: a jammed sample where $L=10.5$ $mm$ in black and an unjammed sample where $L=12.2$ $mm$ in blue. The jammed regime is seen as a region of high tangent modulus at low strain, as indicated by the pink line. The softened regime is indicated with the green line. (c) Schematic of jammed knitted fabric phase diagram. For some possible combinations of yarn length $L$ and yarn radius $r$, the fabric is not knittable, indicated by the grey region. When the fabric is knittable and $L/r$ is smaller than some critical threshold $(L/r)^{*}$, there are some applied stresses that show jammed responses and some that show softened elastic behavior. When $L/r>(L/r)^{*}$, the fabric does not exhibit jammed behavior even for small stresses. We are unable to quantify $(L/r)^{*}$ using our simulation methods, but it is discussed in more detail in Supplementary Section VIII.
@@
When an external force is applied to unjammed knitted fabrics, there is a wide linear elastic regime (extending anywhere from 10% to 300% strain) before the gradual onset of strain-stiffening behavior becomes evident [18, 19, 20, 21, 15]. The linear elastic moduli and strain range are highly dependent on stitch pattering and yarn properties [15, 16]. Knitted fabrics have a slip-knot microstructure that allow the constituent yarn to change in shape, giving rise to a highly extensible response [22, 23]. Because of this, knitted fabrics are ideal for wearable electronics [24], soft robotics [25], and medical applications [26].
@@
However, in highly dense knits (which can be manufactured by machine by applying additional tension to the yarn during knitting or by hand by using needles that are small compared to the yarn radius), a much stiffer linear regime develops for low strain. This low-stress high-rigidity regime then softens into low stiffness behavior at intermediate strain (Fig. 1b). Other work has recorded this behavior in experimental stress-strain data and either attributed it to frictional affects [20] or ignored it completely. Low-strain-high-modulus behavior has been seen in fabrics made of both compressible [27, 28, 15] and incompressible yarns [19, 29, 20]. For fabrics used in technological implementations such as strain sensors [30], a jammed response can limit the sensitivity of the sensor.
@@
In this Letter, we examine the macroscopic response of knitted fabrics, both experimentally and computationally, to determine a classification for jammed, softened, and unjammed regimes as we change the density – as approximated by the dimensionless parameter $L/r$, where the length of yarn per stitch $L$ is normalized by yarn radius $r$ – and the applied stress. We use simulations to determine the micromechanical origins of this jammed behavior, and we determine that in the jammed regime individual stitches are geometrically confined such that the yarn must rearrange in order to accommodate applied stress. In these densely-knitted fabrics, the stitches are highly compressed, which restricts the ability of the yarn to change shape. This results in a stiffer response at low strain. The interplay between yarn bending and the change in contacts between segments of yarn within a stitch leads to anisotropic jamming mechanisms, where geometric changes mediated by bending stiffness cause jamming in the $x$-direction and changing contacts due to compression causes jamming in the $y$-direction. This anisotropic jamming mechanism can be seen in the aspect ratio of the fabric, the energy of individual stitches, and the forces generated by yarn contacts.
@@ II Experimental Observations of Jamming
We fabricated experimental samples of stockinette fabric made from lace-weight yarn on a STOLL industrial knitting machine, using the fabrication and methods outlined in [15] and additionally described in Supplementary Section I. Each sample was made of the same yarn with identical manufacturing parameters except for the “stitch size,” which is a dimensionless parameter within the M1 Plus software that effectively changes the yarn tension of the machine. Changing the yarn tension leads to changes in the length of yarn per stitch and yarn radius within the samples, as described in Supplementary Section II. As the “stitch size” decreases, both the length of yarn per stitch and yarn radius also decrease, as seen in Fig  S2. Images of experimental samples being stretched can be seen in Fig. 2a,b. Stress-strain results for these fabrics can be seen in Fig. S1.
Jamming can be found in experimental samples for multiple different “stitch size” settings (Fig.  S1). As the “stitch size” decreases, the fabrics become tighter (Fig. 5c-d), less elastic, and begin to show jammed behavior (Fig.  S1). All fabrics that display jammed behavior in the $y$-direction are also show it in the $x$-direction, but the converse does not hold. Since both the length of yarn per stitch and yarn radius change in these samples (Fig. S2), one or more of these changes leads to changes in jamming.
@@ III Modeling Micromechanics
To investigate the micro-mechanical details of the jammed regime, we utilized the simulation framework described in [15]. We include a brief overview here. These simulations model a single unit cell of the knitted fabric. For stockinette fabric, the unit cell is a single knit stitch, shown in Fig. 1a. In our model, two main types of deformation to the yarn cost energy: bending and compression. Yarn bending is modeled through elastica theory, such that the total bending energy $E_{b}$ is given by
@@
The compression force experienced when two non-neighboring segments of yarn come into contact is modeled as an inverse power law for a soft-shell hard-core yarn [31, 32, 15, 16] in terms of a non-dimensional measure of distance the yarn is compressed, $\zeta\equiv(R-2r_{\rm core})/(2r-2r_{\rm core})$ where $r$ is the yarn radius, $r_{\rm core}$ is radius of the incompressible inner core, and $R$ is the straight-line separation between two segments of yarn. These quantities are diagrammed in Fig. S3. For $\zeta<1$, the yarn is compressed, and the compression energy is given by:
$E_{c}(\zeta)=k\frac{(2r-2r_{\rm core})^{2}}{p(p-1)}\iint{\rm d}s\,{\rm d}s^{\prime}\left[\zeta^{1-p}(s,s^{\prime})-1-(p-1)\left(1-\zeta(s,s^{\prime})\right)\right]$ (2)
@@
where the compression constant $k$ sets the energy scale, $p$ is the exponent characterizing the nonlinearity of the force law, and the double integral is taken over all yarn segments. Both the compression constant $k$ and characteristic exponent $p$ are fit to experimental compression data as described in [15]. The total yarn energy is thus the sum of bending and compression energies, $E_{\rm total}=E_{b}+E_{c}$.
Figure 2: (a) Photograph of the experimental sample made at the smallest sampled machine tension setting (10.5) before it is stretched by the UTM. (b) Photograph of the same sample at the maximum stretch with an applied force of 60 N. The stress-strain plots all experimental samples can be found in Fig. S1. Photo credit: A. Patrick Cachine. (c) Experimental measurements of the secant modulus $Y_{\rm sec}$ for fabrics of different machine tension settings stretched in the x- (pink) and y-directions (green). (d) Experimental measurements of the offset stress $\sigma_{0}$ for fabrics of different machine tension settings stretched in the x- (pink) and y-directions (green). The large experimental error in the measured $L/r$ values is primarily from error in the measurement of $r$, which is discussed in Supplementary Section II an shown in Fig. S2. (e) Simulation values for the secant modulus $Y_{\rm sec}$ for sweeps of the length of yarn per stitch $L$ (gray), the yarn radius $r$ (orange), and yarn radius for fixed ratio of $r_{\rm core}/r$ (dark blue) stretched in the x-direction. (f) Simulation values for the offset stress $\sigma_{0}$ for sweeps of the length of yarn per stitch $L$ (gray), the yarn radius $r$ (orange), and yarn radius for fixed ratio of $r_{\rm core}/r$ (dark blue) stretched in the x-direction. The yarn radius sweeps had a length of yarn per stitch $L=12.2$ mm. The yarn radius for fixed ratio of $r_{\rm core}/r$ had a length of yarn per stitch $L=12.0$ mm. All other simulation parameters are as described in 1. Results for fabrics stretched in the y-direction are similar and shown in Fig. S4. Equations for the lines of best fit for c-f can be found in Table S1.
@@
The simulation takes an initial configuration of yarn with correct stitch topology and then relaxes its shape to find the minimum energy configuration, while constraining the length of yarn in the simulation box to be fixed. The yarn center-line is modeled as a $\mathcal{C}^{2}$ continuous curve through a Bézier spline representation using degree-5 Bernstein polynomials (see [15] for more details). To generate a stress-strain curve from the simulation, many simulations are run for different unit cell dimensions to represent the fabric being pulled in one direction. Unlike the simulation method in [15], the zero-force configuration is found by running the simulation with only the length constraint. Instead of sweeping over one of the stitch cell dimensions and interpolating to predict the zero-force configuration, both dimensions of the unit cell are allowed to vary freely to find the global minimum energy state of the fabric. From that zero-force state, we then fix the $x$- (or $y$-) dimension of the unit cell, allowing the other dimension to vary freely and the yarn shape to change. We iteratively repeat this process with a larger and larger $x$- (or $y$-) dimension until we reach a stress of 0.4 N/mm. Notably, these simulations solve for a minimum energy and do not incorporate yarn friction. In previous results with a similar simulation framework, these static simulations are able to generate jammed fabric behavior with the characteristic high modulus at low strain (see Figure 2 in [15]). Unless otherwise specified, the simulation inputs are as described in Table 1.

### v2

@@ Abstract
We investigate the mechanical behavior of jammed knitted fabrics, where geometric confinement leads to an initially stiff mechanical response that softens into low stiffness behavior with additional applied stress. We show that the jammed regime is distinguished by changes in yarn geometry and contact rearrangement that must occur to allow the individual stitches to stretch. These rearrangements allow for the relaxation of high residual stresses that are present within jammed fabric, altering the low-strain response. We demonstrate that fabric jamming can be induced or reduced by changing either the constituent yarn or the fabric manufacturing parameters. Analysis of experimental samples shows that changing the “stitch size” in manufacturing affects both the in situ yarn radius and the length of yarn per stitch, both of which affect jamming.
@@ I Introduction
Granular systems are ubiquitous in nature and manufacturing, from sandy beaches to corn-filled silos, and they exhibit some common behaviors. Generally, jamming is the process through which an amorphous or disordered system attains a yield stress through changing properties such as density or temperature [1, 2]. This disorder differentiates jamming from crystallization, which gives rise to an ordered, equilibrium state. For athermal hard spheres, the onset of jamming signifies the sudden transition from zero to non-zero pressure response when the granular system is compressed [2]. A universal feature of granular systems is a jamming transition, where a system which has fluid-like behavior – flowing under constant stress – develops compressional rigidity and behaves like a solid under constant stress.
@@
The transition to, or from, a jammed state often occurs due to changes in constraints on the constituent particles. For hard particles, these constraints arise from contacts between neighbors, and the jamming transition can be rationalized by simple constraint counting and thus characterized by the average number of contacts per particle [3]. By contrast, the equilibrium configurations of “elasto-granular” systems that combine hard particle jamming with fiber elasticity involve non-local, compressive and tensile stress distributions, giving rise to new mechanical behaviors [4, 5]. For moderately deformable systems, like foams, the number of contacts still plays a large role in the characterization of the jamming transition [6, 7]. In highly deformable systems – where the deformation is of the order of the particle size – additional constraints, such as particle shape [8, 9], contribute to the onset of jammed behavior.
Jamming has been introduced as a particularly useful property in mechanical metamaterials and soft robotics, where the variable stiffness of the material leads to novel applications [10, 11, 12] such as soft grippers that conform around an object and utilize jamming to rigidify and move the object [13]. In this vein, we examine jamming in a non-granular context, where the continuous, flexible yarn that is initially confined in tightly-knitted fabrics must rearrange in order to support higher stresses.
@@
Knitting is a fabric manufacturing method that creates some of the most commonly used clothing items, including t-shirts and socks. Knitted fabrics are comprised of a network of interlocked loops formed by a continuous piece of yarn (Fig. 1a). In weft knitting, a single strand of yarn forms a series of loops horizontally within a row. In the next row, a new series of loops are threaded through the loops on the prior row. Each loop of yarn is called a stitch. Combinations of stitches that have different topologies creates to different elastic responses and textures in the final fabric [14, 15]. Fabrics made entirely of knit stitches are called stockinette fabric, or jersey, and is the fabric type found in the body of t-shirts. This is the type of fabric we will study here, and it is the stiffest of the four most common knitted fabric types [15, 16].
@@
The term “jamming,” as applied to knitted textiles, was coined by Postle [17] to describe an anomalous stiff response to low strains. Jamming in knitted fabrics is qualitatively different from jamming in granular systems in a multitude of ways. For one, knitted fabrics are far from amorphous, as they typically possess periodic translational symmetry. Unlike jammed, athermal granular systems, the fabrics show jammed behavior under tension and move from a jammed state to a softened state (Fig. 1) with greater applied force [17]. Here, we argue that the change in mechanical properties due to yarn confinement in tightly-knitted fabrics can be regarded as analogous to the jamming behavior seen in other athermal, granular systems.
@@
Figure 1: (a) Simulated knitted fabric with the unit cell highlighted. This graphic is made from a simulation where $k=0.1~{}{\rm mN/mm^{-2}}$, $r=0.74~{}{\rm mm}$, and $L=13.0~{}{\rm mm}$. However, we rendered the image with a yarn radius $r=0.5~{}{\rm mm}$ instead of the simulation value to better display the yarn structure. (b) Stress versus strain in the x-direction for two different simulated fabrics: a jammed sample where $L=10.5~{}{\rm mm}$ in black and an unjammed sample where $L=12.2~{}{\rm mm}$ in green. The jammed regime is seen as a region of high tangent modulus at low strain, as indicated by the pink line. The softened regime is indicated with the blue line. (c) Schematic of jammed knitted fabric phase diagram. For some possible combinations of yarn length $L$ and yarn radius $r$, the fabric is not knittable, indicated by the grey region. When the fabric is knittable and $L/r$ is smaller than some critical threshold $(L/r)^{*}$, there are some applied stresses that show jammed responses and some that show softened elastic behavior. When $L/r>(L/r)^{*}$, the fabric does not exhibit jammed behavior even for small stresses. We are unable to quantify $(L/r)^{*}$ using our simulation methods, but it is discussed in more detail in Supplementary Section VIII.
@@
When an external force is applied to unjammed knitted fabrics, there is a wide linear elastic regime (extending anywhere from $\sim$10to $\sim$300strain) before the gradual onset of strain-stiffening behavior becomes evident [18, 19, 20, 21, 15]. The linear elastic moduli and strain range are highly dependent on stitch pattering and yarn properties [15, 16]. Knitted fabrics have a slip-knot microstructure that allow the constituent yarn to change in shape, giving rise to a highly extensible response [22, 23] compared to woven, felted, and other non-woven fabrics. Because of this, knitted fabrics are ideal for wearable electronics [24], soft robotics [25], and medical applications [26].
@@
However, in highly dense knits (which can be manufactured by machine by applying additional tension to the yarn during knitting or by hand by using needles that are small compared to the yarn radius), a much stiffer linear regime develops for low strain. This low-stress high-rigidity regime then softens into low stiffness behavior at intermediate strain (Fig. 1b). Other work has recorded this behavior in experimental stress-strain data and either attributed it to frictional effects [20] or ignored it completely. This low-strain-high-modulus behavior has been seen in fabrics made of both compressible [27, 28, 15] and incompressible yarns [19, 29, 20]. For fabrics used in technological implementations such as strain sensors [30], a jammed response can limit the sensitivity of the sensor.
Figure 2: (a) Photograph of an experimental sample made at the smallest sampled “stitch size” setting (10.5) before it is stretched by the UTM. (b) Photograph of the same sample at the maximum stretch with an applied force of 60 N. The stress-strain plots for all experimental samples can be found in Fig. S1. Photo credit: A. Patrick Cachine. (c) Experimental measurements of the secant modulus $Y_{\rm sec}$ for fabrics of different “stitch size” settings stretched in the $x$- (purple) and $y$-directions (blue). (d) Experimental measurements of the offset stress $\sigma_{0}$ for fabrics of different “stitch size” settings stretched in the $x$- and $y$-directions (purple and blue, respectively). The error bars in the experimentally measured $L/r$ values are standard errors of the mean, discussed in Supplementary Section II and shown in Fig. S2. (e) Values for the secant modulus $Y_{\rm sec}$ for simulation sweeps of the length of yarn per stitch $L$ (gray), the yarn radius $r$ (orange), and yarn radius for fixed ratio of $r_{\rm core}/r$ (dark blue) stretched in the $x$-direction. (f) Values for the offset stress $\sigma_{0}$ for simulation sweeps of the length of yarn per stitch $L$ (gray), the yarn radius $r$ (orange), and yarn radius for fixed ratio of $r_{\rm core}/r$ (dark blue) stretched in the $x$-direction. The yarn radius sweeps had a length of yarn per stitch $L=12.2~{}{\rm mm}$. The yarn radius for fixed ratio of $r_{\rm core}/r$ had a length of yarn per stitch $L=12.0~{}{\rm mm}$. All other simulation parameters are as described in 1. Results for fabrics stretched in the y-direction are similar and shown in Fig. S4. Equations for the lines of best fit for c-f can be found in Table S1.
@@
Here, we examine the macroscopic response of knitted fabrics, both experimentally and computationally, to determine a classification for jammed, softened, and unjammed regimes as we change the applied stress and the “density” – as approximated by the dimensionless parameter $L/r$, where the length of yarn per stitch $L$ is normalized by yarn radius $r$. We use simulations to determine the micromechanical origins of this jammed behavior, and we determine that in the jammed regime individual stitches are geometrically confined such that the yarn must rearrange in order to accommodate applied stress. In these densely-knitted fabrics, the stitches are highly compressed, which restricts the ability of the yarn to change shape to accommodate stretching. This results in a stiffer response at low strain. The interplay between yarn bending and the change in contacts between segments of yarn within a stitch leads to anisotropic jamming mechanisms, where geometric changes mediated by bending stiffness cause jamming in the $x$-direction and changing contacts due to compression causes jamming in the $y$-direction. This anisotropic jamming mechanism can be seen in the aspect ratio of the fabric, the energy of individual stitches, and the forces generated by yarn contacts.
@@ II Experimental Observations of Jamming
We fabricated experimental samples of stockinette fabric made from lace-weight yarn on a STOLL industrial knitting machine, using the fabrication and methods outlined in [15] and additionally described in Supplementary Section I. Each sample was made of the same yarn with identical manufacturing parameters except for the “stitch size,” which is a dimensionless parameter within the M1 Plus software that effectively changes the yarn tension of the machine. Changing the yarn tension leads to changes in the length of yarn per stitch and yarn radius within the samples, as described in Supplementary Section II. Both the length of yarn per stitch and yarn radius decrease as the “stitch size” parameter decreases, as seen in Fig  S2. Images of experimental samples being stretched can be seen in Fig. 2a,b. Stress-strain results for these fabrics can be seen in Fig. S1.
Jamming can be found in experimental samples for multiple different “stitch size” settings (Fig. S1). As the “stitch size” decreases, the fabrics become tighter (Fig. 5c-d), stiffer, and begin to show jammed behavior (Fig. S1). All fabrics that display jammed behavior in the $y$-direction also show it in the $x$-direction, but the converse does not hold. Since both the length of yarn per stitch and yarn radius change in these samples (Fig. S2), one or more of these changes leads to changes in jamming.
@@ III Modeling Micromechanics
To investigate the micro-mechanical details of the jammed regime, we utilized the simulation framework described in [15]. We include a brief overview here. These simulations model a single unit cell of the knitted fabric. For stockinette fabric, the unit cell is a single knit stitch, shown in Fig. 1a. In our model, two main types of deformation to the yarn cost energy: bending and compression. Yarn bending is modeled as Euler elastica, such that the total bending energy $E_{b}$ is given by
@@
The compression force experienced when two non-neighboring segments of yarn come into contact is modeled as an inverse power law for a soft-shell, hard-core yarn [31, 32, 15, 16] in terms of a non-dimensional measure of the distance that the yarn is compressed, $\zeta(s,s^{\prime})\equiv\left(R(s,s^{\prime})-2r_{\rm core}\right)/(2r-2r_{\rm core}),$ where $r$ is the yarn radius, $r_{\rm core}$ is radius of the incompressible inner core, and $R(s,s^{\prime})$ is the straight-line separation between two segments of yarn that are parameterized by arclengths $s$ and $s^{\prime}$, respectively. These quantities are diagrammed in Fig. S3. For $\zeta<1$, the yarn is compressed, and the compression energy is given by:
$E_{c}(\zeta)=k\frac{(2r-2r_{\rm core})^{2}}{p(p-1)}\iint{\rm d}s\,{\rm d}s^{\prime}\left[\zeta^{1-p}(s,s^{\prime})-1-(p-1)\left(1-\zeta(s,s^{\prime})\right)\right]\,\,,$ (2)
@@
where the compression constant $k$ sets the energy scale, $p$ is the exponent characterizing the nonlinearity of the force law, and the double integral is taken over all yarn segments. Both the compression constant $k$ and characteristic exponent $p$ are fit to experimental compression data [15]. The total yarn energy is thus the sum of bending and compression energies, $E_{\rm total}=E_{b}+E_{c}$.
@@
The simulation takes an initial configuration of yarn with correct stitch topology and then relaxes its shape to find the minimum energy configuration, while constraining the length of yarn in the simulation box to be fixed. The yarn center-line is modeled as a $\mathcal{C}^{2}$ continuous curve through a Bézier spline representation using degree-5 Bernstein polynomials (see [15] for more details). Unless otherwise specified, the simulation inputs are as described in Table 1. To generate a stress-strain curve from the simulation, many simulations are run for different unit cell dimensions to represent the fabric being pulled in one direction. Unlike the simulation method in [15], the zero-force configuration is found by running the simulation with only the length constraint. Instead of sweeping over one of the stitch cell dimensions and interpolating to predict the zero-force configuration, both dimensions of the unit cell are allowed to vary freely to find the global minimum energy state of the fabric. From that zero-force state, we then fix the $x$- (or $y$-) dimension of the unit cell, allowing the other dimension to vary freely and the yarn shape to change. We iteratively repeat this process with a larger and larger $x$- (or $y$-) dimension until we reach a stress of 0.4 N/mm. Notably, these simulations solve for a minimum energy and do not incorporate yarn friction. In previous results with a similar simulation framework, these static simulations are able to generate jammed fabric behavior with the characteristic high modulus at low strain (see Figure 2 in [15]).